calculate the molarity of kmno4 by using a cbr of a 6ml of 0.1m
1
Expert's answer
2021-07-14T05:05:17-0400

Molarity = number of moles/Volume

Molar Mass of CBr = 331.627

= 331.627 × 0.1 = 33.1627

= 6/33.1627 = 0.181 M

Molar Mass of 158.034

= 0.181/0.158.034

= 1.15m/L